OpenStudy (anonymous):

I always got confused with these problems. can anyone help? find the y coordinate of the point on the graph of 6X+3Y+1=0 with an x coordinate of -3

OpenStudy (nincompoop):

what part is confusing?

OpenStudy (rational):

you don't need to graph this equation to find the y coordinate on the graph

OpenStudy (rational):

although it would be good see how this simple equation represents all the infinity points on entire line

OpenStudy (anonymous):

I guess the 0 is what throws me off

OpenStudy (nincompoop):

oh that is easy laughing out loud try this one, put the format Ax + By + c = 0 into Ax + By +c - c = 0 - c

OpenStudy (nincompoop):

it is just natural that there is a zero at the end of the equation if the value of c is on the left hand side standard form would be Ax + By = C

OpenStudy (nincompoop):

so what that really means is that your standard equation of a line is really 6x + 3y = -1

OpenStudy (rational):

actually 0 is a good number, you should not get pissed off so much seeing it on the right hand side :P

OpenStudy (nincompoop):

respond accordingly @KarolinaJosephine

OpenStudy (anonymous):

what?

OpenStudy (nincompoop):

at least give us a feedback that you are understanding what I am trying to communicate

OpenStudy (anonymous):

sorry my internet is screwing up and keeps saying connection lost so I have to keep refreshing it. a storm is blowing through my area. and no I still don't understand how the apparent answer is 17/3

OpenStudy (phi):

start with 6X+3Y+1=0 with an x coordinate of -3 that means replace the x (i.e erase the x and put -3 in its place) 6(-3) + 3y+1 =0 remember 6x is short for 6 times x, or in this case 6 times -3 6*-3 is -18 -18 + 3y + 1 =0 do you know how to simplify -18 + 1 ?

OpenStudy (anonymous):

but why is it positive instead of negative?

OpenStudy (phi):

do you know how to simplify -18 + 1 ?

OpenStudy (phi):

you should get -18+1 = -17 that makes -18 + 3y + 1 =0 3y -17 = 0

OpenStudy (phi):

the next step is to add +17 to both sides and simplify 3y -17 +17 = 0+17 can you simplify the left side? can you simplify the right side? what do you get?

OpenStudy (anonymous):

I'm not 5 ok.

OpenStudy (nincompoop):

laughing out loud

OpenStudy (nincompoop):

@abb0t

OpenStudy (phi):

If you don't type in the answer I have no idea if you are following or not.

OpenStudy (phi):

If you know how to simplify 3y -17 +17 = 0+17 then post it. If you don't know, then say you don't know.

OpenStudy (abb0t):

i dont thnk i get this high in class. mrs johnson only on chaptr 3. do u go 2 connexus???